Calculate the pH of a 9.1 x 10 Maqueous solution of hydroiodic acid (HI). Round your answer to 2 decimal places. 5 ?
Answer #1

Answer! HI Strong acid a [HI] = 9.1x108m = [ht] <107 so [H+J = 9,1x168 +1o7 from water = 10740.91 + 1) [H+J = 1914107 pH = -l